### v2

# Bounds on decaying sterile neutrinos via magnetic dipole moment from COB intensity

Abstract
A recent observation by Long Range Reconnaissance Imager (LORRI) mounted on NASA’s New Horizons yielded the most accurate measurement of the cosmic optical background (COB). The reported COB intensity is $11.16\pm 1.35$ $\mathrm{nW/m^{2}/sr}$ at a pivot wavelength $\lambda_{piv}=0.608\,\mu\mathrm{m}$ observed in the range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$. After subtracting the measured intensity from the deep Hubble Space Telescope count, diffused galactic light, and scattered light from bright star foregrounds, an anomalous intensity $2.99\pm 2.03\leavevmode\nobreak\ \mathrm{nW/m^{2}/sr}$ has been found. We considered radiatively decaying sterile neutrinos of mass scales keV and eV, as dark matter candidate and as dark radiation respectively, that could contribute to this anomalous reported intensity. Using this, we derive upper bounds on sterile-to-sterile and active-to-sterile transition magnetic dipole moments. We find that sterile neutrinos with mass of $\mathcal{O}(\rm keV)$ scale take values of the transition magnetic moment in the range $3\times 10^{-12}\,\mu_{B}-10^{-9}\,\mu_{B}$, whereas for mass of $\mathcal{O}(\rm eV)$ scale the transition magnetic moment in the range $3\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$, here $\mu_{B}$ is Bohr magneton.
Sterile-to-sterile transition magnetic dipole moment, Sterile-to-active transition magnetic dipole moment, Low Energy Effective Field Theory(LEFT)
I Introduction
Line-intensity mapping (LIM) is an observational method for extragalactic astronomy and cosmology that measures the integrated intensity of spectral lines emitted from galaxies and the intergalactic medium at a specific observed frequency Kovetz et al. (2017, 2020). Since LIM experiments capture information from all incoming photons, they have the potential to detect electromagnetic radiation produced by dark matter decays directly. For instance, in Ref. Grin et al. (2007), authors have derived a stringent upper bound on the axion-photon coupling strength by searching for optical line emission produced from decaying axions. In Ref. Gong et al. (2016), authors have shown that photons produced from decaying axions with eV mass can explain the anisotropy of near-IR extragalactic background light. In Ref. Creque-Sarbinowski and Kamionkowski (2018), authors have proposed that LIM experiments can detect photons produced from decaying and annihilating dark matter particles. However, measuring the extragalactic background light in the optical band is a difficult task. The primary reason is mitigating the overwhelming foreground, which requires accurate modelling and well-calibrated instruments Bernal et al. (2022). In Ref. Zemcov et al. (2017a), authors have suggested observing such radiations in the optical spectrum—the cosmic optical background (COB) is possible if one can eliminate the foreground radiations, primarily solar in nature.
The Long Range Reconnaissance Imager (LORRI) mounted on NASA New Horizons spacecraft, currently operating at a distance $51.3$ astronomical unit (AU) from the Sun, has been measuring the COB photon intensity in the wavelength range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$ $(\sim 1.3\leavevmode\nobreak\ \rm eV-3\leavevmode\nobreak\ \,\rm eV)$ Zemcov et al. (2017b); Lauer et al. (2021, 2022); Symons et al. (2023); Postman et al. (2024). In $2021$, LORRI operating at 42-45 AU from the Sun reported an optical intensity from unknown sources of $8.8\pm 4.9\,\rm nW/m^{2}/sr$ $11.9\pm 4.6\,\rm nW/m^{2}/sr$ Lauer et al. (2021). Later, it yielded an anomalous intensity of $8.06\pm 1.92\,\rm nW/m^{2}/sr$ while operating at $51.3$ AU Lauer et al. (2022). In Lauer et al. (2022), the authors considered diffused galactic light (DGL) from the Milky Way and scattered light from bright stars (SSL) contributing to the total sky level. Further, they subtracted the intensity from integrated light of external galaxies (IGL), known as deep Hubble galaxy count, of $8.17\pm 1.18\,\rm nW/m^{2}/sr$ from the total COB level to report the anomalous intensity Lauer et al. (2022). Later, this result was also supported by an independent research following blind methodology Symons et al. (2023). However, in a recent analysis Postman et al. (2024), the authors have reworked the analysis of the foregrounds from the DGL and found the contribution of DGL to the COB intensity is greater compared to the one found in article Lauer et al. (2022). Thus yielding an anomalous intensity of $2.99\pm 2.03\,\rm{nW/m^{2}/sr}$ in the optical spectrum, relatively modest to the IGL intensity Postman et al. (2024).
This anomalous photon excess may indicate the presence of physics beyond the Standard Model. Various studies have explored possible explanations, such as photons emitted from decaying axions with masses in the range of $\sim 5\leavevmode\nobreak\ \rm eV-25\leavevmode\nobreak\ \rm eV$ Bernal et al. (2022); Nakayama and Yin (2022); Carenza et al. (2023). Additionally, decaying sterile neutrinos could produce a similar photon signal through their decay into photons and neutrinos.
Short-baseline neutrino experiments have hinted at the existence of additional neutrinos, and numerous studies have investigated sterile neutrinos as a potential dark matter candidate Ho and Scherrer (2013). The investigation of photon radiated from the sterile neutrinos can provide an exciting probe for the beyond standard model containing sterile neutrinos.
Motivated by these possibilities, we analyze photon flux originated from decaying sterile neutrinos. We also explore it as source of the observed anomalous intensity.
[…]
In our work, we focus on the keV mass scale for sterile neutrinos, as they are considered a viable dark matter candidate, as well as we have considered the eV mass scale of decaying sterile neutrinos. Rather than considering a specific model, we adopt an effective field theory approach Beltrán et al. (2024), which makes our analysis more general and less dependent on particular model assumptions. Anomalies observed in short-baseline neutrino experiments, such as MiniBooNE Aguilar-Arevalo et al. (2021, 2018) and LSND Aguilar et al. (2001), have sparked significant interest in the possibility of eV-scale sterile neutrinos, which have been extensively studied in the literature. Motivated by these scenarios, we investigate the potential role of both the keV- and eV-scale sterile neutrinos in contributing to the reported anomalous COB intensity, thus deriving upper bounds on sterile-to-sterile and sterile-to-active transition magnetic dipole moment. Future experiments like the SPHEREx telescope surveying the sky at wavelengths between $0.75-5\,\mu m$ might shed light on the properties of radiatively decaying sterile neutrinos Crill et al. (2020); Doré et al. (2014, 2016, 2018).
The paper is organised as follows. In Section II, we discuss the radiative decay process of sterile neutrinos through the sterile-to-sterile and sterile-to-active transition magnetic dipole moment in the low-energy effective field theory. In Section III, we derive the mean specific intensity for the keV mass range of decaying sterile neutrinos. In Section IV, we have analyzed our results by determining the minimal required magnetic dipole moment and decay width for sterile-to-sterile and active-to-sterile transition magnetic dipole moment and the existing limits on it. In this section, we have considered the two COB measurements. Here, we also derive the novel constraints on sterile-to-sterile transition magnetic dipole moment through the contribution on the COB from optical to X-rays region. In Section V, we present our conclusions with an outlook for future works.
II Radiative Decay of Sterile Neutrino through Transition Magnetic Dipole Moment
We consider Standard model augmented with sterile neutrinos in the effective field theory approach. Assuming sterile neutrino mass scales well below the electroweak scale, we can write the effective Lagrangian contributes to the magnetic moment of the neutrino sector:
[…]
where, $d_{\rm NN\gamma}$ represents the sterile-to-sterile transition magnetic dipole moment and $d_{\rm\nu N\gamma}$ denotes the active-to-sterile transition magnetic dipole moment, respectively. The term $\mathcal{O}_{\rm NN\gamma}^{ij}=(\bar{N}_{Ri}^{c}\sigma_{\mu\nu}N_{Rj})F^{\mu\nu}$, $\mathcal{O}_{\nu N\gamma}^{\alpha i}=(\bar{\nu}_{L\alpha}^{c}\sigma_{\mu\nu}N_{Ri})F^{\mu\nu}$ represent the effective field operators, where $F^{\mu\nu}$ is the electromagnetic field strength tensor. The electric and magnetic field off-diagonal non-zero dipole moment components give rise to radiative decay through two different sterile states. Therefore, in this radiative decay process ($m_{1}\rightarrow m_{2}+\gamma$), the decay width of sterile neutrino can be written as Atre et al. (2009); Bondarenko et al. (2018); Balantekin and Vassh (2014)
$\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{2|d_{\rm NN\gamma}|^{2}}{\pi}m_{1}^{3}\left(1-\frac{m_{2}^{2}}{m_{1}^{2}}\right)^{3}.$ (2)
By defining $\delta=1-\frac{m_{2}}{m_{1}}$, which represents the mass splitting between two eigenstates of sterile neutrino, we can rewrite Eq. (2) as
$\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{2|d_{\rm NN\gamma}|^{2}}{\pi}m_{1}^{3}\left(2-\delta\right)^{3}\delta^{3},$ (3)
From Eq. (3), we can derive the sterile-to-sterile transition magnetic dipole moment $d_{\rm NN\gamma}$.
There are several experimental upper bounds on the neutrino magnetic moment. Recent studies of the NOMAND neutrino detector at CERN put constraints on the active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim(10^{-6}-10^{-9})\mu_{B}$ Gninenko and Krasnikov (1999). Here, $\mu_{B}\sim 3\times 10^{-7}\,\rm eV^{-1}$ represents Bohr magneton. In an anti-neutrino electron scattering experiment, by examining the electron recoil spectra, it was found that active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 10^{-9}\mu_{B}$ Derbin et al. (1993). The Super-Kamiokande experiment put an upper bound limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.1\times 10^{-10}\mu_{B}$ Liu et al. (2004), and the Borexino collaboration put the upper bound limit at $90\%$ confidence level as $5.4\times 10^{-11}\mu_{B}$ Agostini et al. (2017); Arpesella et al. (2008). MUNU collaboration has found the upper bound on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 9.0\times 10^{-11}\mu_{B}$ Babu et al. (2020), whereas TEXONO collaboration has obtained this active-to-sterile transition magnetic dipole moment upper bound limit $d_{\rm\nu N\gamma}\lesssim 7.4\times 10^{-11}\mu_{B}$ Balantekin and Vassh (2014). The Planck+BAO put the constraint on the active-to-sterile transition magnetic dipole moment, and it varies from $3.7\times 10^{-11}\mu_{B}$ to $9.1\times 10^{-12}\mu_{B}$ Carenza et al. (2024). Recent studies analysing the cooling of red giant stars have found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Capozzi and Raffelt (2020). The anomalous stellar cooling due to plasmon decay has been found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}$ is $d_{\rm\nu N\gamma}\lesssim 3\times 10^{-12}\mu_{B}$ Alexander (2016).
III Mean specific intensity of decaying sterile neutrinos
Consider a radiative decaying particle of rest mass $m_{\chi}$ represented as $\chi\to\gamma+\xi$, where $\gamma$ and $\xi$ represent a photon and another particle, respectively. The wavelength of the photon produced in the rest frame of $\chi$ is Chen and Kamionkowski (2004); Bernal et al. (2021)
$\lambda_{e}=\frac{hc}{m_{\chi}c^{2}}\times\frac{2}{\left(1-m_{\xi}^{2}/m_{\chi}^{2}\right)}$ (4)
where $h$ and $c$ are the Planck’s constant and the speed of light in vacuum, respectively. We define the second term on the right-hand side of Eq. (4) as $x^{-1}=\frac{2}{\left(1-m_{\xi}^{2}/m_{\chi}^{2}\right)}$. We note that for a two-photon decay process, $x=1/2$. The wavelength of photon produced from a particle decaying at redshift $z$ will be observed today as $\lambda_{obs}=(1+z)\lambda_{e}$. If these photons are not absorbed by the intergalactic gas along the light of sight, then they will be observed today as the extra-galactic background lights. The redshift dependence of the energy density of decaying dark matter particles is given by $\rho_{\chi}(t)=\rho_{\chi}(0)(1+z)^{3}\mathrm{e}^{-\Gamma_{\chi}t_{\rm U}}$, where $t_{\rm U}\sim 10^{18}\,\rm sec$ is the age of the universe, $\Gamma_{\chi}$ represents the decay rate, and $\rho_{\chi}(0)$ is the dark matter energy density today. If $\Gamma_{\chi}$ of decaying dark matter particle is such that $\Gamma_{\chi}t_{\rm U}\lesssim 10^{-2}$, we can then consider $e^{-\Gamma_{\chi}t_{\rm U}}\approx 1$. Later in this section, we have shown that the decay rate considered in this work justifies this approximation. The specific intensity of the photons $I_{\nu}$ observed at frequency $\nu_{\rm obs}$ today is given by Creque-Sarbinowski and Kamionkowski (2018)
[…]
where $\epsilon_{\nu_{\rm obs}}$ is the specific luminosity density and $H(z)$ is the Hubble parameter. For a Dirac delta decaying profile, $\epsilon_{\nu_{\rm obs}}$ is expressed as Chen and Kamionkowski (2004); Creque-Sarbinowski and Kamionkowski (2018)
$\epsilon_{\nu_{\rm obs}}(z)=h\mathrm{N}_{\gamma}\Gamma_{\chi}\mathrm{n}_{\chi}(z)\mathrm{E}_{\gamma}\,\delta(\rm E-E_{\gamma}),$ (6)
where, $\rm N_{\gamma}$ and $\rm n_{\chi}(z)$ represent the number of photons produced and the number density of the decaying particles. On solving Eq. (5) for the aforementioned $\epsilon_{\nu_{\rm obs}}$ and using the relation, $I_{\lambda}=I_{\nu}\left|\mathrm{d\nu}/\mathrm{d\lambda}\right|$ we calculate the mean specific intensity per observed wavelength as
$I_{\lambda}=\frac{c}{4\pi}\frac{\Omega_{\chi,0}\rho_{c}\Gamma_{\chi}}{\lambda_{\rm obs}(1+z)\text{H}(z)m_{\chi}}\leavevmode\nobreak\ \frac{hc}{\lambda_{e}}N_{\gamma},$ (7)
where $\Omega_{\chi,0}$ represents the present-day dark matter density parameter and $\rho_{c}$ is the critical density. In this work, we consider a decaying sterile neutrino of mass $m_{1}$ to produce a photon and another sterile neutrino of mass $m_{2}$ comparable to the mass of $m_{1}$. In this case, the term $hc/\lambda_{e}$ in Eq. (7) can be replaced with $x\,m_{1}c^{2}$. Considering both $m_{1}$ and $m_{2}$ are of $\mathcal{O}(\rm keV)$, we can approximate $m_{1}+m_{2}\sim 2m_{1}$. The photon’s energy produced from this process becomes $E_{\gamma}=m_{1}-m_{2}$. From hereinafter, we define $m_{1}-m_{2}=\Delta m$. Consequently, the term $x$ mentioned earlier will be $\Delta m/m_{1}$. On substituting all the terms as mentioned above in Eq. (7), we rewrite the equation as
$I_{\lambda}=\frac{c}{4\pi}\frac{{\Omega_{\chi,0}}{\rho_{c}}{c^{2}}\Gamma_{m_{1}}}{\lambda_{obs}(1+z)H(z)}\leavevmode\nobreak\ \frac{\Delta m}{m_{1}}.$ (8)
From the above equation, we can observe that $I_{\lambda}$ is directly proportional to the decay rate and the energy of the released photon. However, it is inversely related to the mass of the decaying particle and decays down over redshifts due to the expansion of the Universe.
IV Results and Discussions
In the previous section, we discussed that the intensity of photon production by the sterile neutrino radiative decays which is linked to the neutrino magnetic moment. Here, we investigate the reported COB anomalous intensities, that is $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al. (2022) and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ Postman et al. (2024), and analyze them within the framework of the neutrino magnetic moment to derive upper bounds on the mass and magnetic moment. From hereinafter we will represent the reported intensities $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ as $I_{806}$ Lauer et al. (2022) and $I_{299}$ respectively Postman et al. (2024), unless stated otherwise. Specifically, we investigate three possible scenarios: sterile-sterile, sterile-active, and active-active neutrino magnetic moments.
IV.1 Sterile to sterile neutrino decay
Figure 1: (a) Illustrates the variation in specific intensity $I_{\lambda}$ with respect to $\lambda_{\rm obs}$, for $m_{1}=1\,\rm keV$ and $3\,\rm keV$ decaying at $\Gamma_{m_{1}}=10^{-22}\,\text{sec}^{-1}$ and $\Gamma_{m_{1}}=10^{-21}\,\text{sec}^{-1}$ respectively— depicted in the blue solid and red dashed lines . (b) The variations in $I_{\lambda}$ contributing to LORRI’s reported anomalous intensities. For $I=2.99\pm 2.03\,\rm nW/m^{2}/sr$ Postman et al. (2024), we consider $1-40\,\rm keV$ sterile neutrinos decaying in the range $\Gamma_{m_{1}}\sim(10^{-22}-10^{-21})\,\text{sec}^{-1}$ – shown in blue solid lines. Further for $I=8.06\pm 1.92\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ Lauer et al. (2022), we consider $1-20\,\rm keV$ sterile neutrinos decaying in the range $\Gamma_{m_{1}}\sim(10^{-21}-10^{-20})\,\text{sec}^{-1}$ – shown in red dashed lines. The green-shaded area illustrates the shape of LORRI’s responsivity for observed wavelengths Bernal et al. (2022).
Figure 2: Top panel: The minimum required decay width to obtain $I_{806}$ and $I_{299}$ shown in red dashed and blue solid lines, respectively, for masses $1\leavevmode\nobreak\ \rm keV-40\leavevmode\nobreak\ \rm keV$. Bottom panel: The corresponding upper bounds on $d_{\rm NN\gamma}$ for the anomalous intensities $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024).
Figure 3: This figure illustrates the upper bounds on $d_{\rm NN\gamma}$ versus the mass of the sterile neutrino $(m_{1})$. The red dashed and blue solid line denote $d_{NN\gamma}$ obtained for intensity $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024). The lavender-shaded region represents the excluded values of $d_{\rm NN\gamma}-m_{1}$ values from X-ray observations Natwariya and Nayak (2022).
In this section, we calculate the required decay rate and the transition magnetic moment for the anomalous intensities reported by LORRI Lauer et al. (2022); Postman et al. (2024). We consider the keV mass range sterile neutrino as a dark matter candidate, which decays into another sterile neutrino and photon through the radiative decay process. We note that the sterile neutrinos with mass $<0.4\,\rm keV$ are excluded as a dark matter candidate by a conservative application of Tremaine-Gunn bound Abazajian and Koushiappas (2006). Now, to calculate $I_{\lambda}$ required to explain the anomalous intensity of COB, we need to determine the energy of photons $(\Delta m)$, as shown in Eq. (8). The LORRI observation reported flux of photons at pivot wavelength $0.608\,\mu m\,(\sim 2\,\rm eV)$ Lauer et al. (2022). Therefore, $\Delta m$ should be in the range $2\,\rm eV-10\,\rm eV$, such that LORRI’s band could observe these redshifting photons today. We restrict the energies of photons in this range, as the photons produced at $z\gtrsim 4$ will get absorbed by the intergalactic gas medium Madau (1995); Inoue et al. (2014).
To illustrate the variation of $I_{\lambda}$ with $\lambda_{\rm obs}$, we considered $1\,\rm keV$ and $3\,\rm keV$ sterile neutrinos decaying at $10^{-22}\,\rm sec^{-1}$ and $10^{-21}\,\rm sec^{-1}$, respectively— shown in blue solid lines and red dashed lines, respectively, in Fig. 1. In this figure, the cutoff vertical line for wavelength $\lambda_{e}$ is the wavelength of photons observed today. The green-shaded region represents the shape responsivity of LORRI’s camera with respect to the observed wavelength Weaver et al. (2020); Bernal et al. (2022); Porras-Bedmar et al. (2024). Further, we fixed the intensities to the LORRI’s reported values, shown in Fig. 1. We first fix the obtained intensity to $I_{806}$ Lauer et al. (2022), for $1-20\,\rm keV$ sterile neutrinos decaying in the range $\Gamma_{m_{1}}\sim\left(10^{-21}-10^{-20}\right)\,\rm sec^{-1}$, as depicted in the red dashed lines. Similarly, the blue solid lines show $1-40\,\rm keV$ sterile neutrinos decaying at $\Gamma_{m_{1}}\sim\left(10^{-22}-10^{-21}\right)\,\rm sec^{-1}$ producing the LORRI’s reported intensity $I_{299}$ Postman et al. (2024). We note that the intensity curves at lower $\lambda_{\rm obs}$ values suggests production photons with energy $\Delta m$ for the aforementioned $m_{1}$ and $\Gamma_{m_{1}}$ originating from higher redshifts. From Eq. (8), we can observe that high energetic photons originating at higher redshifts require smaller $\Gamma_{m_{1}}$. Therefore, a sterile neutrino with a large decay width is required to produce less energetic photons in order to contribute to the anomalous COB intensities. In the next section, we present the minimum required $\Gamma_{m_{1}}$, and hence the $d_{\rm NN\gamma}$, to obtain the reported COB anomalous intensities for different $m_{1}$ values.
In the top panel of Fig. (2), we plot the required minimum decay width for different masses to obtain the anomalous intensities. The red dashed and blue solid lines represent the minimum decay width for masses $1-40\,\rm keV$ in order to obtain the reported intensities in $I_{806}$ and $I_{299}$, respectively. We find that sterile neutrinos of mass $3\,\rm keV$ and $1\,\rm keV$ require $1\times 10^{-21}\,\rm sec^{-1}$ and $8.54\times 10^{-22}\,\rm sec^{-1}$ for intensities reported in $I_{8.06\pm 1.92}$ and $I_{2.99\pm 2.03}$, respectively. The decay width of the sterile neutrino increases as we increase the mass. This can be analysed from the fact that $I_{\lambda}\propto\Gamma_{m_{1}}/m_{1}$— shown in Eq. (8). We restrict our analysis to $1\leq m_{1}/\rm keV\leq 40$ because a further increase in $m_{1}$ requires an increase in $\Gamma_{m_{1}}$. For instance, the required $\Gamma_{m_{1}}$ for $m_{1}\gtrsim 40\,\rm keV$ requires $\gtrsim 10^{-20}\,\rm sec^{-1}$ for $I_{806}$, which become comparable to $\sim t^{-1}_{\rm U}$ which requires a consideration of variation in the dark matter density— as mentioned in section (III). We then plot the minimum values of $d_{\rm NN\gamma}$, in the bottom panel, for the above obtained $\Gamma_{m_{1}}$ values — shown in the red dashed and blue solid lines. We find that under the conditions $m_{1}+m_{2}\sim 2m_{1}$ and $m_{1}-m_{2}=\Delta m$, the decay width Eq. 2 becomes $16|d_{\rm NN\gamma}|^{2}\,\Delta m^{3}/\pi$. As mentioned earlier, $\Delta m$ takes values of $2\,\rm eV-10\,\rm eV$ for the observed flux of photons. Therefore, to calculate upper bounds on $d_{\rm NN\gamma}$, we consider the minimum values of $\Gamma_{m_{1}}$ for the corresponding $m_{1}$ values while fixing $\Delta m=10\,\rm eV$. We find that the transition magnetic moment $d_{\rm NN\gamma}/\rm eV^{-1}$ takes values of $(3\times 10^{-11}-10^{-8})\mu_{B}$ and $(3\times 10^{-12}-10^{-9})\mu_{B}$ for mass $1\,\rm keV-40\,\rm keV$— as shown in the red dashed and blue solid lines, respectively. Further in Fig. (3), we present the upper bounds on $d_{\rm NN\gamma}$ in the presence of existing X-ray bounds on $\Gamma_{m_{1}}-m_{1}$ values as the lavender coloured region Natwariya and Nayak (2022). In the next section, we derive constraints on the sterile-to-active transition magnetic dipole moment $(d_{\nu N\gamma})$ by considering the COB anomalous intensities $I_{806}$ and $I_{299}$.
IV.2 Sterile to active neutrino decay
In this section, we consider the $\mathcal{O}$(eV) mass scale of sterile neutrino, which is decaying into active neutrino and photon via active-to-sterile transition magnetic dipole moment. This radiative decay process can be expressed as $N_{1}\to\nu_{a}+\gamma$. The wavelength of the emitted photon $(\lambda_{e})$ can be calculated from Eq. (4). The term $\lambda_{e}$ depends on both the mass of sterile and active neutrino. Here, we fix the active neutrino mass to be $0.12\,\rm eV$. As explained earlier in section (IV.1), the required energy of the photons should be between $2\,\rm eV-10\,\rm eV$ such that these redshifting photons will be observed today as an excess radiation in the COB. Using these conditions, we find that the mass of the sterile neutrino will be in the range of $4\,\rm eV-20\,\rm eV$. The upper bound on the mass of the sterile neutrino to be a dark matter candidate, derived from the Traemaine-Gunn bound, is $0.4$ keV Abazajian and Koushiappas (2006). Therefore, we examine the $\mathcal{O}$(eV) scale decaying sterile neutrinos, which may be considered as dark radiation. In the standard scenario case, only three neutrino species contribute to the $\rm N_{eff}^{SM}$ which is $3.044$. However, in a nonstandard scenario, the effective neutrino species can be written as $\rm N_{eff}=N_{eff}^{SM}+\Delta N_{eff}$, where $\Delta\rm N_{eff}$ comes from the additional relativistic components that take part in the dark radiation Abazajian and Koushiappas (2006). Dark radiation measures the amount of radiation energy contributed by the relativistic species except photons. Usually, neutrino oscillation anomalies give the idea for searching the new additional relativistic particles in the cosmos Abazajian and Koushiappas (2006). Motivated by the neutrino oscillation experiment for explaining the dark radiation, the mass range of sterile neutrino should be $m_{s}\lesssim\rm eV$ Drewes (2013). So here, we take the mass range $(4\,\rm eV-20\,\rm eV)$ of decaying sterile neutrino, which can be regarded as dark radiation.
Figure 4: (a): Represent the variations in specific intensity $I_{\lambda}$ with respect to $\lambda_{\rm obs}$, for $12$ eV sterile neutrinos decaying at $10^{-23}\,\text{sec}^{-1}$ and $10^{-22}\,\text{sec}^{-1}$, shown in blue solid and red dashed lines, respectively. (b): Shows variations in $I_{\lambda}$ for $12$ eV sterile neutrinos of masses with decay rates $10^{-23}-10^{-22}\,\rm sec^{-1}$ for fixed intensities $I_{806}$ and $I_{299}$, as shown in the red dashed and blue solid lines, respectively. The green-shaded area illustrates the responsivity curve of LORRI for each detected wavelength.
The decay width for a decaying sterile neutrino via sterile-to-active transition magnetic moment is given by Beltrán et al. (2024)
$\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{|d_{\rm\nu N\gamma}|^{2}}{2\pi}m_{1}^{3}\,,$ (9)
where $d_{\nu\rm N\gamma}$, $m_{1}$, and $m_{2}$ represent the active-to-sterile transition magnetic dipole moment and the mass of the sterile and active neutrino, respectively. On the contrary to the decay width shown in Eq. (2), here $\Gamma_{m_{1}}$ is independent of the mass difference between the sterile and active neutrino. We then write the expression for the mean specific intensity $I_{\lambda}$ of photons produced from eV-scale decaying sterile neutrino as
$I_{\lambda}=x\,\frac{c}{4\pi}\frac{{\Omega_{s}}{\rho_{c}}{c^{2}}\Gamma_{m_{1}}}{\lambda_{obs}(1+z)H(z)}.$ (10)
Here, the term $x=1/2\,\left(1-m_{2}^{2}/m_{1}^{2}\right)$ is similar to the one shown in Eq. (4). Whereas, $\Omega_{s}$ represents the density parameter of sterile neutrinos, which can be expressed as $\Omega_{s}h^{2}=m_{1}/94.1\,\rm eV$ for a nonthermal scenario, which comes from the “Dodelson-Widrow Mechanism” Dodelson and Widrow (1994).
Using Eq. (10), we calculate $I_{\lambda}$ for $(4\,\rm eV-20\,\rm eV)$ mass range of sterile neutrinos, that can contribute to the reported intensities by LORRI in the wavelength range of $(0.4\,\rm\mu m-0.9\,\rm\mu m)$. For a fixed $m_{1}=12\,\rm eV$, we considered decay widths $\Gamma_{m_{1}}$ of $10^{-23}\,\rm sec^{-1}$ and $-10^{-22}\,\rm sec^{-1}$ to plot variations in $I_{\lambda}$ for different $\lambda_{\rm obs}$ values— shown in blue solid and red dashed lines, respectively, in Fig. 4. Further in Fig. 4, we vary $m_{1}$ in the range $(4\,\rm eV-20\,\rm eV)$ and $\Gamma_{m_{1}}$ within $10^{-23}-10^{-22}\,\rm sec^{-1}$, while fixing the total intensity of photons to $I_{806}$ and $I_{299}$— as shown in red dashed and blue solid lines, respectively. Below, we calculate the decay rate for mass $m_{1}$ required to obtain both the COB intensities and derive upper bounds on sterile-to-sterile transition magnetic dipole moment.
Figure 5: Top panel: The minimum required decay rate to put constraint on sterile-to-active transition magnetic dipole moment ($d_{\nu N\gamma}$) by taking into consideration the contribution of the anomalous intensity of COB $2.99\pm 2.03\,\rm nW/m^{2}/sr$ Postman et al. (2024) and the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al. (2022) reported by LORRI for sterile neutrinos with masses ($4\leavevmode\nobreak\ \rm eV-20\leavevmode\nobreak\ \rm eV)$. Bottom panel: The corresponding minimum magnetic coupling strength $(d_{\rm\nu N\gamma})$ in $\text{eV}^{-1}$.
In the top panel of Fig. 5, the blue solid and red dashed lines show the minimum decay width $\Gamma_{m_{1}}$ for sterile neutrino $4\leavevmode\nobreak\ \text{eV}\leq m_{1}\leq 20\leavevmode\nobreak\ \text{eV}$ to obtain the total intensities of $I_{299}$ and $I_{806}$, respectively. We find that $\Gamma_{m_{1}}$ increases for heavier sterile neutrinos. This can be analyzed from the fact that the specific intensity $(I_{\lambda})$ depends directly on the decay width while mass enters only through redshift $z(\lambda_{\rm obs},m_{1})$ which can only alter the Hubble parameter $H(z)$— as shown in Eq. (10). Thus, larger mass increases $z(\lambda_{\rm obs},m_{1})$, suggesting an early decay of sterile neutrino. However, as the intensity of photons redshifts more for an early decay, a larger decay width is required to obtain the required intensity today. Further, the solid blue and dashed red lines in the bottom panel show upper bounds on active-to-sterile transition magnetic moment $d_{\nu N\gamma}$ derived for the above blue solid line and red dashed line respectively using Eq. (9). We observe that $d_{\nu N\gamma}$ decreases with the increase of the mass of the decaying sterile neutrino. This can be analysed from Eq. (9), where $d_{\nu N\gamma}$ is inversely proportional to the mass of decaying sterile neutrinos. We find that, the required $d_{\nu N\gamma}$ to explain the anomalous intensities of $I_{806}$ and $I_{299}$ lies between $3.33\times 10^{-12}\,\rm\mu_{B}-10^{-11}\,\rm\mu_{B}$ and $3.33\times 10^{-13}\,\rm\mu_{B}-10^{-12}\,\rm\mu_{B}$, respectively, for sterile neutrino mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$.
Currently, the upper bound on the active-to-sterile transition dipole moment $d_{\rm\nu N\gamma}$ for sterile neutrino’s mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$ remains unconstrained. There are several experimental upper bounds on the active-to-sterile transition magnetic dipole moment. Recent studies of XENON1T found that the upper bound on $d_{\nu N\gamma}\lesssim 6\times 10^{-11}\mu_{B}$ for several MeV mass range of sterile neutrino Shoemaker et al. (2021). However, the recent analysis of global clusters sets an upper limit as $d_{\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Cañas et al. (2016).
IV.3 Active-to-active neutrino magnetic moment
In the UV model, the active-to-active neutrino magnetic moment is naturally suppressed Beltrán et al. (2024). Active-sterile mixing can create an active-to-active neutrino magnetic moment in the broken phase parameters when the right-handed neutrinos are integrated out Beltrán et al. (2024). As such $d_{\nu\nu\gamma}\lesssim 2\times 10^{-8}\leavevmode\nobreak\ \text{\rm GeV}^{-1}$, we may anticipate that the active-to-active neutrino magnetic moments can satisfy the constraints from TEXONO Li et al. (2003), GEMMA Beda et al. (2010), LSND, and Borexino Agostini et al. (2017) in the inverse see-saw mechanism, where larger active-sterile mixing can occur. For the specified mass range $(0.12\leavevmode\nobreak\ \text{\rm eV}-0.7\leavevmode\nobreak\ \text{\rm eV})$ of active neutrinos, we are not able to obtain the specific intensity of photons due to the radiative decay of active neutrinos. As for this aforementioned mass range of active neutrinos, a very less number of photons can be generated, which cannot explain the COB anomalous intensity in the observed wavelength range. For that reason, we cannot observe any specific intensity of photons in the observed wavelength range $\lambda_{obs}=0.2\leavevmode\nobreak\ \mu m-1.2\leavevmode\nobreak\ \mu m$ because to explain the COB anomaly, we need minimum $4$ eV mass of sterile neutrinos. To summarize, we put constraints on the sterile-to-sterile and sterile-to-active transition magnetic dipole moment for ($1\,\rm keV-20\,\rm keV$) and ($4\,\rm eV-20\,\rm eV$) mass range of stile neutrinos, respectively by considering two COB measurements.
V Conclusion and Outlook
Recent measurements of the cosmic optical background (COB) photon flux intensity by LORRI have revealed some excess intensity with no known sources. This excess may indicate a potential source of energetic photons beyond the Standard Model (BSM). In this study, we considered the possibility of decaying sterile neutrinos producing photons and explored the corresponding decay rates. Notably, we did not account for photon energy generation through cascades involving other decay products. We found the required decay rate of sterile neutrino and obtained upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments of the order $\mathcal{O}(10^{-22}-10^{-21})\leavevmode\nobreak\ \text{sec}^{-1}$ for the keV mass range of the decaying sterile neutrino and for the $\mathcal{O}$(eV) mass range of the decaying sterile neutrino, the decay rate should be in the range of $\mathcal{O}(10^{-23}-10^{-22})\leavevmode\nobreak\ \text{sec}^{-1}$. Here, we adopt a minimal framework involving sterile neutrinos using the effective field theory approach. The radiative decay of sterile neutrinos, resulting in photon production, is primarily governed by magnetic moments associated with sterile-to-sterile $(d_{NN\gamma})$ and sterile-to-active neutrino ($d_{\nu N\gamma}$) transitions. We calculate the $d_{NN\gamma}$ and $d_{\nu N\gamma}$ for various mass scales and decay rates. We find that, $d_{NN\gamma}$ take values of $3\times 10^{-11}\,\mu_{B}-10^{-9}\,\mu_{B}$ for sterile neutrinos of mass range $1\,\rm keV-40\,\rm keV$. Whereas sterile neutrinos, with mass range $4\,\rm eV-20\,\rm eV$, decaying via sterile-to-active transition magnetic moment $(d_{\nu\rm N\gamma})$ take values of $3.33\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$ for newly found COB measurement. In our work, we explain the bounds on sterile-to-sterile and sterile-to-active transition magnetic moments by considering the two COB measurements: one is $2.99\pm 2.03\,\rm nW/m^{2}/sr$ intensity, and another one is due to anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$. Here, we have plotted the specific intensity for both COB measurements. Motivated by this work, one can probe CUB, CXB, and CIB measurements to constrain the sterile-to-sterile and sterile-to-active transition magnetic dipole moment. We leave this as future work. Some upcoming experiments, such as GALEX and ULTRASAT satellites, can probe and constrain the decaying dark matter in the ultraviolet bandwidth Libanore and Kovetz (2025). Our minimal framework based on decaying sterile neutrinos of mass $\mathcal{O}$(keV) and $\mathcal{O}$(eV) can contribute to the unknown excess COB intensity, thus deriving upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments.
VI Acknowledgement
We sincerely acknowledge the usage of python packages scipy [^1] https://www.scipy.org/ and astropy [^2] https://www.astropy.org/index.html. We thank Rahul Kothari for his earlier collaboration on this project.

### v3

@@
-Bounds on decaying sterile neutrinos via magnetic dipole moment from COB intensity
+compat=1.1.0
 Abstract
@@
 The reported COB intensity is $11.16\pm 1.35$ $\mathrm{nW/m^{2}/sr}$ at a pivot wavelength $\lambda_{piv}=0.608\,\mu\mathrm{m}$ observed in the range $0.4\,\mu\mathrm{m}\lesssim\lambda\lesssim 0.9\,\mu\mathrm{m}$.
-After subtracting the measured intensity from the deep Hubble Space Telescope count, diffused galactic light, and scattered light from bright star foregrounds, an anomalous intensity $2.99\pm 2.03\leavevmode\nobreak\ \mathrm{nW/m^{2}/sr}$ has been found.
-We considered radiatively decaying sterile neutrinos of mass scales keV and eV, as dark matter candidate and as dark radiation respectively, that could contribute to this anomalous reported intensity.
-Using this, we derive upper bounds on sterile-to-sterile and active-to-sterile transition magnetic dipole moments.
-We find that sterile neutrinos with mass of $\mathcal{O}(\rm keV)$ scale take values of the transition magnetic moment in the range $3\times 10^{-12}\,\mu_{B}-10^{-9}\,\mu_{B}$, whereas for mass of $\mathcal{O}(\rm eV)$ scale the transition magnetic moment in the range $3\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$, here $\mu_{B}$ is Bohr magneton.
-Sterile-to-sterile transition magnetic dipole moment, Sterile-to-active transition magnetic dipole moment, Low Energy Effective Field Theory(LEFT)
+After subtracting the measured intensity from the deep Hubble Space Telescope count, diffused galactic light, and scattered light from bright star foregrounds, an anomalous intensity of $2.99\pm 2.03\leavevmode\nobreak\ \mathrm{nW/m^{2}/sr}$ has been found.
+We considered radiatively decaying sterile neutrinos of keV mass scale, as dark matter candidate, that could contribute to this anomalous reported intensity.
+Using this, we derive upper bounds on the sterile-to-sterile transition magnetic moment.
+We find that sterile neutrinos with mass of $\mathcal{O}(\rm keV)$ scale take values of the transition magnetic moment in the range $3\times 10^{-13}\,\rm eV^{-1}-10^{-9}\,\rm eV^{-1}$ to explain the anomalous intensity of $2.99\pm 2.03\,\rm nW/m^{2}/sr$.
+Sterile-to-sterile transition magnetic moment, Low Energy Effective Field Theory (LEFT)
 I Introduction
@@
 In Ref.
-Creque-Sarbinowski and Kamionkowski (2018), authors have proposed that LIM experiments can detect photons produced from decaying and annihilating dark matter particles.
+Creque-Sarbinowski and Kamionkowski (2018), the authors have proposed that LIM experiments can detect photons produced from decaying and annihilating dark matter particles.
 However, measuring the extragalactic background light in the optical band is a difficult task.
@@
 (2024).
-In $2021$, LORRI operating at 42-45 AU from the Sun reported an optical intensity from unknown sources of $8.8\pm 4.9\,\rm nW/m^{2}/sr$ $11.9\pm 4.6\,\rm nW/m^{2}/sr$ Lauer et al.
+In $2021$, LORRI operating at 42-45 AU from the Sun reported an optical intensity from unknown sources of $8.8\pm 4.9\,\rm nW/m^{2}/sr$ - $11.9\pm 4.6\,\rm nW/m^{2}/sr$ Lauer et al.
 (2021).
-Later, it yielded an anomalous intensity of $8.06\pm 1.92\,\rm nW/m^{2}/sr$ while operating at $51.3$ AU Lauer et al.
+Later, it yielded an anomalous intensity of COB, $8.06\pm 1.92\,\rm nW/m^{2}/sr$ while operating at $51.3$ AU Lauer et al.
 (2022).
-In Lauer et al.
+In Ref.
+Lauer et al.
 (2022), the authors considered diffused galactic light (DGL) from the Milky Way and scattered light from bright stars (SSL) contributing to the total sky level.
-Further, they subtracted the intensity from integrated light of external galaxies (IGL), known as deep Hubble galaxy count, of $8.17\pm 1.18\,\rm nW/m^{2}/sr$ from the total COB level to report the anomalous intensity Lauer et al.
+Further, they subtracted the intensity from integrated light of external galaxies (IGL), known as deep Hubble galaxy count, $8.17\pm 1.18\,\rm nW/m^{2}/sr$ from the total COB level to report the anomalous intensity Lauer et al.
 (2022).
@@
 However, in a recent analysis Postman et al.
-(2024), the authors have reworked the analysis of the foregrounds from the DGL and found the contribution of DGL to the COB intensity is greater compared to the one found in article Lauer et al.
+(2024), the authors have reworked the analysis of the foregrounds from the DGL and found that the contribution of DGL to the COB intensity is greater compared to the one found in the article Lauer et al.
 (2022).
-Thus yielding an anomalous intensity of $2.99\pm 2.03\,\rm{nW/m^{2}/sr}$ in the optical spectrum, relatively modest to the IGL intensity Postman et al.
+Thus yielding an anomalous intensity $2.99\pm 2.03\,\rm{nW/m^{2}/sr}$ in the optical spectrum, relatively modest to the IGL intensity Postman et al.
 (2024).
 This anomalous photon excess may indicate the presence of physics beyond the Standard Model.
-Various studies have explored possible explanations, such as photons emitted from decaying axions with masses in the range of $\sim 5\leavevmode\nobreak\ \rm eV-25\leavevmode\nobreak\ \rm eV$ Bernal et al.
+Various studies have explored such possibilities, for e.g., photons emitted from decaying axions with masses in the range of $\sim 5\leavevmode\nobreak\ \rm eV-25\leavevmode\nobreak\ \rm eV$ Bernal et al.
 (2022); Nakayama and Yin (2022); Carenza et al.
@@
 The investigation of photon radiated from the sterile neutrinos can provide an exciting probe for the beyond standard model containing sterile neutrinos.
-Motivated by these possibilities, we analyze photon flux originated from decaying sterile neutrinos. We also explore it as source of the observed anomalous intensity.
+Motivated by these possibilities, we analyse photon flux originating from decaying sterile neutrinos.
+We also explore it as the source of the observed anomalous intensity.
 […]
-In our work, we focus on the keV mass scale for sterile neutrinos, as they are considered a viable dark matter candidate, as well as we have considered the eV mass scale of decaying sterile neutrinos.
+In this work, we focus on the keV mass scale of sterile neutrinos, as they are considered a viable dark matter candidate.
 Rather than considering a specific model, we adopt an effective field theory approach Beltrán et al.
 (2024), which makes our analysis more general and less dependent on particular model assumptions.
-Anomalies observed in short-baseline neutrino experiments, such as MiniBooNE Aguilar-Arevalo et al.
-(2021, 2018) and LSND Aguilar et al.
-(2001), have sparked significant interest in the possibility of eV-scale sterile neutrinos, which have been extensively studied in the literature.
-Motivated by these scenarios, we investigate the potential role of both the keV- and eV-scale sterile neutrinos in contributing to the reported anomalous COB intensity, thus deriving upper bounds on sterile-to-sterile and sterile-to-active transition magnetic dipole moment.
-Future experiments like the SPHEREx telescope surveying the sky at wavelengths between $0.75-5\,\mu m$ might shed light on the properties of radiatively decaying sterile neutrinos Crill et al.
+We investigate the potential role of the keV-mass scale sterile neutrinos in contributing to the reported anomalous COB intensity, thus deriving upper bounds on the sterile-to-sterile transition magnetic moment by varying the $\Delta m$ values from $2\,\rm eV$ to $10\,\rm eV$.
+Future experiments like the SPHEREx telescope surveying the sky at wavelengths $0.75-5\,\mu m$ might shed light on the properties of radiatively decaying sterile neutrinos Crill et al.
 (2020); Doré et al.
@@
 The paper is organised as follows.
-In Section II, we discuss the radiative decay process of sterile neutrinos through the sterile-to-sterile and sterile-to-active transition magnetic dipole moment in the low-energy effective field theory.
+In Section II, we discuss the radiative decay process of sterile neutrinos through the sterile-to-sterile transition magnetic dipole moment in the low-energy effective field theory.
 In Section III, we derive the mean specific intensity for the keV mass range of decaying sterile neutrinos.
-In Section IV, we have analyzed our results by determining the minimal required magnetic dipole moment and decay width for sterile-to-sterile and active-to-sterile transition magnetic dipole moment and the existing limits on it.
-In this section, we have considered the two COB measurements.
-Here, we also derive the novel constraints on sterile-to-sterile transition magnetic dipole moment through the contribution on the COB from optical to X-rays region.
-In Section V, we present our conclusions with an outlook for future works.
-II Radiative Decay of Sterile Neutrino through Transition Magnetic Dipole Moment
+In Section IV, we have analyzed our results by determining the minimal required sterile-to-sterile transition magnetic moment and decay width of keV-mass scale sterile neutrinos.
+In this section, we consider the two COB measurements; one is anomalous COB excess intensity, $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and another one is the rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$.
+Here, we also derive the novel constraints on the sterile-to-sterile transition magnetic moment through the contribution of the COB intesnity from the optical to X-rays region.
+In sectionV, we explain the UV origin of the small mass splitting of sterile neutrinos.
+In Section VI, we present our conclusions with an outlook for future works.
+II Radiative Decay of Sterile Neutrino through Transition Magnetic Moment
 We consider Standard Model augmented with sterile neutrinos in the effective field theory approach.
-Assuming sterile neutrino mass scales well below the electroweak scale, we can write the effective Lagrangian contributes to the magnetic moment of the neutrino sector:
-[…]
-where, $d_{\rm NN\gamma}$ represents the sterile-to-sterile transition magnetic dipole moment and $d_{\rm\nu N\gamma}$ denotes the active-to-sterile transition magnetic dipole moment, respectively.
+Assuming sterile neutrino mass scales well below the electroweak scale, we can write the effective Lagrangian density that contributes to the magnetic moment of the neutrino sector:
+$\mathcal{L}_{N_{R}LEFT}\supset d_{\rm NN\gamma}^{ij}\mathcal{O}_{\rm NN\gamma}^{ij}+d_{\nu N\gamma}^{\beta i}\mathcal{O}_{\nu N\gamma}^{\beta i}+\text{h.c.}\,$, (1)
+where, $d_{\rm NN\gamma}$ represents the sterile-to-sterile transition magnetic moment and $d_{\rm\nu N\gamma}$ denotes the active-to-sterile transition magnetic moment, respectively.
 The term $\mathcal{O}_{\rm NN\gamma}^{ij}=(\bar{N}_{Ri}^{c}\sigma_{\mu\nu}N_{Rj})F^{\mu\nu}$, $\mathcal{O}_{\nu N\gamma}^{\alpha i}=(\bar{\nu}_{L\alpha}^{c}\sigma_{\mu\nu}N_{Ri})F^{\mu\nu}$ represent the effective field operators, where $F^{\mu\nu}$ is the electromagnetic field strength tensor.
-The electric and magnetic field off-diagonal non-zero dipole moment components give rise to radiative decay through two different sterile states.
+The electric and magnetic field off-diagonal non-zero transition dipole moment components give rise to radiative decay through two different sterile states.
 Therefore, in this radiative decay process ($m_{1}\rightarrow m_{2}+\gamma$), the decay width of sterile neutrino can be written as Atre et al.
@@
 $\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{2|d_{\rm NN\gamma}|^{2}}{\pi}m_{1}^{3}\left(2-\delta\right)^{3}\delta^{3},$ (3)
-From Eq. (3), we can derive the sterile-to-sterile transition magnetic dipole moment $d_{\rm NN\gamma}$.
+From Eq. (3), we can derive the sterile-to-sterile transition magnetic moment $d_{\rm NN\gamma}$.
 There are several experimental upper bounds on the neutrino magnetic moment.
-Recent studies of the NOMAND neutrino detector at CERN put constraints on the active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim(10^{-6}-10^{-9})\mu_{B}$ Gninenko and Krasnikov (1999).
-Here, $\mu_{B}\sim 3\times 10^{-7}\,\rm eV^{-1}$ represents Bohr magneton.
-In an anti-neutrino electron scattering experiment, by examining the electron recoil spectra, it was found that active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 10^{-9}\mu_{B}$ Derbin et al.
+Recent studies of the NOMAND neutrino detector at CERN put constraints on the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim(10^{-6}-10^{-9})\mu_{B}$ Gninenko and Krasnikov (1999).
+Here, $\mu_{B}\sim 3\times 10^{-7}\,\rm eV^{-1}$ represents the Bohr magneton.
+In an anti-neutrino electron scattering experiment, by examining the electron recoil spectra, it was found that the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 10^{-9}\mu_{B}$ Derbin et al.
 (1993).
-The Super-Kamiokande experiment put an upper bound limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.1\times 10^{-10}\mu_{B}$ Liu et al.
+The Super-Kamiokande experiment put an upper bound limit on active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 1.1\times 10^{-10}\mu_{B}$ Liu et al.
 (2004), and the Borexino collaboration put the upper bound limit at $90\%$ confidence level as $5.4\times 10^{-11}\mu_{B}$ Agostini et al.
@@
 (2008).
-MUNU collaboration has found the upper bound on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 9.0\times 10^{-11}\mu_{B}$ Babu et al.
-(2020), whereas TEXONO collaboration has obtained this active-to-sterile transition magnetic dipole moment upper bound limit $d_{\rm\nu N\gamma}\lesssim 7.4\times 10^{-11}\mu_{B}$ Balantekin and Vassh (2014).
-The Planck+BAO put the constraint on the active-to-sterile transition magnetic dipole moment, and it varies from $3.7\times 10^{-11}\mu_{B}$ to $9.1\times 10^{-12}\mu_{B}$ Carenza et al.
+MUNU collaboration has found the upper bound on active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}\lesssim 9.0\times 10^{-11}\mu_{B}$ Babu et al.
+(2020), whereas TEXONO collaboration has obtained this active-to-sterile transition magnetic moment upper bound limit $d_{\rm\nu N\gamma}\lesssim 7.4\times 10^{-11}\mu_{B}$ Balantekin and Vassh (2014).
+The Planck+BAO put the constraint on the active-to-sterile transition magnetic moment, and it varies from $3.7\times 10^{-11}\mu_{B}$ to $9.1\times 10^{-12}\mu_{B}$ Carenza et al.
 (2024).
-Recent studies analysing the cooling of red giant stars have found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Capozzi and Raffelt (2020).
-The anomalous stellar cooling due to plasmon decay has been found an upper limit on active-to-sterile transition magnetic dipole moment $d_{\rm\nu N\gamma}$ is $d_{\rm\nu N\gamma}\lesssim 3\times 10^{-12}\mu_{B}$ Alexander (2016).
+Recent studies analysing the cooling of red giant stars have found an upper limit on active-to-sterile transition magnetic moments $d_{\rm\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Capozzi and Raffelt (2020).
+The anomalous stellar cooling due to plasmon decay has been found an upper limit on the active-to-sterile transition magnetic moment $d_{\rm\nu N\gamma}$ is $d_{\rm\nu N\gamma}\lesssim 3\times 10^{-12}\mu_{B}$ Alexander (2016).
 III Mean specific intensity of decaying sterile neutrinos
@@
 $\lambda_{e}=\frac{hc}{m_{\chi}c^{2}}\times\frac{2}{\left(1-m_{\xi}^{2}/m_{\chi}^{2}\right)}$ (4)
-where $h$ and $c$ are the Planck’s constant and the speed of light in vacuum, respectively.
+where $h$ and $c$ are Planck’s constant and the speed of light in vacuum, respectively.
 We define the second term on the right-hand side of Eq.
@@
 The wavelength of photon produced from a particle decaying at redshift $z$ will be observed today as $\lambda_{obs}=(1+z)\lambda_{e}$.
-If these photons are not absorbed by the intergalactic gas along the light of sight, then they will be observed today as the extra-galactic background lights.
-The redshift dependence of the energy density of decaying dark matter particles is given by $\rho_{\chi}(t)=\rho_{\chi}(0)(1+z)^{3}\mathrm{e}^{-\Gamma_{\chi}t_{\rm U}}$, where $t_{\rm U}\sim 10^{18}\,\rm sec$ is the age of the universe, $\Gamma_{\chi}$ represents the decay rate, and $\rho_{\chi}(0)$ is the dark matter energy density today.
-If $\Gamma_{\chi}$ of decaying dark matter particle is such that $\Gamma_{\chi}t_{\rm U}\lesssim 10^{-2}$, we can then consider $e^{-\Gamma_{\chi}t_{\rm U}}\approx 1$.
-Later in this section, we have shown that the decay rate considered in this work justifies this approximation.
+If the intergalactic gas does not absorb these photons along the line of sight, then they will be observed today as the extragalactic background light.
+The redshift dependence of the energy density of decaying dark matter particles is given by $\rho_{\chi}(t)=\rho_{\chi}(0)(1+z)^{3}\mathrm{e}^{-\Gamma_{\chi}t}$, where $\Gamma_{\chi}$ represents the decay width, $t$ is the time variable, and $\rho_{\chi}(0)$ is the dark matter energy density today.
 The specific intensity of the photons $I_{\nu}$ observed at frequency $\nu_{\rm obs}$ today is given by Creque-Sarbinowski and Kamionkowski (2018)
-[…]
-where $\epsilon_{\nu_{\rm obs}}$ is the specific luminosity density and $H(z)$ is the Hubble parameter.
+$I_{\nu}=\dfrac{1}{4\pi}\int_{0}^{\infty}{dz}\dfrac{c}{H(z)}\dfrac{\epsilon_{\nu_{\rm obs}}(z)}{(1+z)^{4}},$ (5)
+where $\epsilon_{\nu_{\rm obs}}$ is the specific luminosity density, and $H(z)$ is the Hubble parameter.
 For a Dirac delta decaying profile, $\epsilon_{\nu_{\rm obs}}$ is expressed as Chen and Kamionkowski (2004); Creque-Sarbinowski and Kamionkowski (2018)
@@
 $I_{\lambda}=\frac{c}{4\pi}\frac{{\Omega_{\chi,0}}{\rho_{c}}{c^{2}}\Gamma_{m_{1}}}{\lambda_{obs}(1+z)H(z)}\leavevmode\nobreak\ \frac{\Delta m}{m_{1}}.$ (8)
-From the above equation, we can observe that $I_{\lambda}$ is directly proportional to the decay rate and the energy of the released photon.
-However, it is inversely related to the mass of the decaying particle and decays down over redshifts due to the expansion of the Universe.
+From the above equation, we can observe that $I_{\lambda}$ is directly proportional to the decay width and the energy of the released photon.
+However, it is inversely related to the mass of the decaying particle and decays over redshifts due to the expansion of the universe.
 IV Results and Discussions
@@
 (2022) and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ Postman et al.
-(2024), and analyze them within the framework of the neutrino magnetic moment to derive upper bounds on the mass and magnetic moment.
+(2024), and analyse them within the framework of the neutrino magnetic moment to derive an upper bound on the transition magnetic moment.
 From hereinafter we will represent the reported intensities $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ as $I_{806}$ Lauer et al.
-(2022) and $I_{299}$ respectively Postman et al.
+(2022) and $I_{299}$ Postman et al.
 (2024), unless stated otherwise.
-Specifically, we investigate three possible scenarios: sterile-sterile, sterile-active, and active-active neutrino magnetic moments.
+Specifically, we investigate the possible scenario: the sterile-to-sterile transition magnetic moment.
 IV.1 Sterile to sterile neutrino decay
@@
 (2024).
-Figure 3: This figure illustrates the upper bounds on $d_{\rm NN\gamma}$ versus the mass of the sterile neutrino $(m_{1})$.
-The red dashed and blue solid line denote $d_{NN\gamma}$ obtained for intensity $I_{806}$ Lauer et al.
+The colour band represents the change of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ in $\rm eV^{-1}$ versus the mass of sterile neutrino $(m_{1})$ in keV scale, where, $\Delta m$ vary from $2-10\,\rm eV$.
+The green one represents the variation of $d_{NN\gamma}$ due to the anomalous rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$.
+Whereas, the yellow one represents the same variation of $d_{NN\gamma}$ due to the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$.
+Figure 3: The yellow and green colour bands illustrate the variation of $d_{\rm NN\gamma}$ versus the mass of the sterile neutrino $(m_{1})$ with the changes of $\Delta m$, which vary from $2\,\rm eV$ to $10\,\rm eV$ .
+These two-colour bands denote the changes of $d_{NN\gamma}$ obtained from anomalous intensities of COB, $I_{806}$ Lauer et al.
 (2022) and $I_{299}$ Postman et al.
-(2024).
-The lavender-shaded region represents the excluded values of $d_{\rm NN\gamma}-m_{1}$ values from X-ray observations Natwariya and Nayak (2022).
-In this section, we calculate the required decay rate and the transition magnetic moment for the anomalous intensities reported by LORRI Lauer et al.
+(2024), respectively.
+The cyan-shaded region represents the excluded values of $d_{\rm NN\gamma}-m_{1}$ values from X-ray observations Natwariya and Nayak (2022).
+In this section, we calculate the required decay width and the transition magnetic moment for the anomalous intensities reported by LORRI Lauer et al.
 (2022); Postman et al.
@@
 We consider the keV mass range sterile neutrino as a dark matter candidate, which decays into another sterile neutrino and photon through the radiative decay process.
-We note that the sterile neutrinos with mass $<0.4\,\rm keV$ are excluded as a dark matter candidate by a conservative application of Tremaine-Gunn bound Abazajian and Koushiappas (2006).
+We note that the sterile neutrinos with mass $0.4\,\rm keV$ are excluded as a dark matter candidate by a conservative application of the Tremaine-Gunn bound Abazajian and Koushiappas (2006).
 Now, to calculate $I_{\lambda}$ required to explain the anomalous intensity of COB, we need to determine the energy of photons $(\Delta m)$, as shown in Eq.
@@
 (2014).
-To illustrate the variation of $I_{\lambda}$ with $\lambda_{\rm obs}$, we considered $1\,\rm keV$ and $3\,\rm keV$ sterile neutrinos decaying at $10^{-22}\,\rm sec^{-1}$ and $10^{-21}\,\rm sec^{-1}$, respectively— shown in blue solid lines and red dashed lines, respectively, in Fig.
+To illustrate the variation of $I_{\lambda}$ with $\lambda_{\rm obs}$, here we consider $1\,\rm keV$ and $3\,\rm keV$ masses of sterile neutrinos decaying at $10^{-22}\,\rm sec^{-1}$ and $10^{-21}\,\rm sec^{-1}$, respectively— shown in blue solid lines and red dashed lines, respectively, in Fig.
 1.
@@
 (2024).
-Further, we fixed the intensities to the LORRI’s reported values, shown in Fig.
+Furthermore, we adjust the intensities to match LORRI’s reported values, as shown in Fig.
 1.
@@
 (2024).
-We note that the intensity curves at lower $\lambda_{\rm obs}$ values suggests production photons with energy $\Delta m$ for the aforementioned $m_{1}$ and $\Gamma_{m_{1}}$ originating from higher redshifts.
+We note that the intensity curves at lower $\lambda_{\rm obs}$ values suggest the production of photons with energy $\Delta m$ for the aforementioned $m_{1}$ and $\Gamma_{m_{1}}$ originating from higher redshifts.
 From Eq.
@@
 In the top panel of Fig.
-(2), we plot the required minimum decay width for different masses to obtain the anomalous intensities.
-The red dashed and blue solid lines represent the minimum decay width for masses $1-40\,\rm keV$ in order to obtain the reported intensities in $I_{806}$ and $I_{299}$, respectively.
+(2), we plot the required minimum decay width for the mass range of sterile neutrinos $1\,\rm keV-40\,\rm keV$ to obtain the anomalous intensities of COB.
+The red dashed and blue solid lines represent the minimum decay width for masses $1-40\,\rm keV$ to obtain the reported intensities in $I_{806}$ and $I_{299}$, respectively.
 We find that sterile neutrinos of mass $3\,\rm keV$ and $1\,\rm keV$ require $1\times 10^{-21}\,\rm sec^{-1}$ and $8.54\times 10^{-22}\,\rm sec^{-1}$ for intensities reported in $I_{8.06\pm 1.92}$ and $I_{2.99\pm 2.03}$, respectively.
@@
 We restrict our analysis to $1\leq m_{1}/\rm keV\leq 40$ because a further increase in $m_{1}$ requires an increase in $\Gamma_{m_{1}}$.
-For instance, the required $\Gamma_{m_{1}}$ for $m_{1}\gtrsim 40\,\rm keV$ requires $\gtrsim 10^{-20}\,\rm sec^{-1}$ for $I_{806}$, which become comparable to $\sim t^{-1}_{\rm U}$ which requires a consideration of variation in the dark matter density— as mentioned in section (III).
+For instance, the required minimum decay width $\Gamma_{m_{1}}$ for $m_{1}\gtrsim 40\,\rm keV$ requires $\gtrsim 10^{-20}\,\rm sec^{-1}$ for $I_{806}$, which becomes comparable to $\sim t^{-1}_{\rm U}$ which requires a consideration of variation in the dark matter density— as mentioned in section (III).
 We then plot the minimum values of $d_{\rm NN\gamma}$, in the bottom panel, for the above obtained $\Gamma_{m_{1}}$ values — shown in the red dashed and blue solid lines.
@@
 As mentioned earlier, $\Delta m$ takes values of $2\,\rm eV-10\,\rm eV$ for the observed flux of photons.
-Therefore, to calculate upper bounds on $d_{\rm NN\gamma}$, we consider the minimum values of $\Gamma_{m_{1}}$ for the corresponding $m_{1}$ values while fixing $\Delta m=10\,\rm eV$.
-We find that the transition magnetic moment $d_{\rm NN\gamma}/\rm eV^{-1}$ takes values of $(3\times 10^{-11}-10^{-8})\mu_{B}$ and $(3\times 10^{-12}-10^{-9})\mu_{B}$ for mass $1\,\rm keV-40\,\rm keV$— as shown in the red dashed and blue solid lines, respectively.
+We find that the sterile-to-sterile transition magnetic moment $d_{\rm NN\gamma}/\rm eV^{-1}$ takes values of $(3\times 10^{-12}-10^{-8})$ and $(3\times 10^{-13}-10^{-9})$ for mass $1\,\rm keV-40\,\rm keV$— as shown in the red dashed and blue solid lines, respectively.
+In Fig.
+2, we illustrate the colour band of $\Delta m$ variation from $2\,\rm eV$ to $10\,\rm eV$ for the two anomalous intensities of COB.
+This colour band represents the changes of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ in $\rm eV^{-1}$ versus the mass of the sterile neutrino $(m_{1})$ for the changes of the minimum required decay width of the sterile neutrino.
+The green colour band represents the changes of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ for the anomalous rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$.
+Whereas, the yellow colour band region represents the changes $d_{NN\gamma}$ for the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$.
+From the decay width Eq.
+(3), we can see that the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ depends directly on the decay width of the sterile neutrino.
+We observe that for the smaller mass difference, $\Delta m=2\leavevmode\nobreak\ \,\rm eV$, the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ approaches up to $10^{-13}\,\rm eV^{-1}$ and the minimum required decay width $(\Gamma_{m_{1}})$ reaches up to $10^{-21}\,\rm sec^{-1}$ for the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$ – as shown in the red dashed line.
+Whereas, for the larger mass difference between two heavy sterile states, $\Delta m=10\leavevmode\nobreak\ \,\rm eV$, the minimum required decay width of the sterile neutrino $(\Gamma_{m_{1}})$ increases and reaches up to $\sim 10^{-19}\,\rm sec^{-1}$ for the mass range of the sterile neutrino $m_{1}\geq 40\,\rm keV$.
+Similarly, the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ also increases and approaches up to $10^{-8}\,\rm eV^{-1}$.
+So, we observe that for the mass range of the sterile neutrino $m_{1}\geq 40\,\rm keV$ or the larger mass difference between two heavy sterile states $\Delta m=10\,\rm eV$, we obtain that the minimum required decay width of the sterile neutrino approaches up to $t_{\,\rm U}\sim 10^{18}\,\rm sec$.
+We illustrate the changes $d_{NN\gamma}$ due to the changes of mass difference $\Delta m$ from $2\,\rm eV$ to $10\,\rm eV$ in this Fig.
+2.
 Further in Fig.
-(3), we present the upper bounds on $d_{\rm NN\gamma}$ in the presence of existing X-ray bounds on $\Gamma_{m_{1}}-m_{1}$ values as the lavender coloured region Natwariya and Nayak (2022).
-In the next section, we derive constraints on the sterile-to-active transition magnetic dipole moment $(d_{\nu N\gamma})$ by considering the COB anomalous intensities $I_{806}$ and $I_{299}$.
+(3), we present the upper bound on the $d_{\rm NN\gamma}$ in the presence of existing X-ray bounds on $\Gamma_{m_{1}}-m_{1}$ values as the cyan-coloured region Natwariya and Nayak (2022).
+For instance, the upper bound on the sterile-to-sterile transition magnetic moment $d_{NN\gamma}$ from X-ray observations is approximately $10^{-6}\,\rm eV^{-1}$ for the mass of sterile neutrino $5\,\rm keV$.
+In contrast, for the larger value of sterile neutrino mass, $40\,\rm keV$, the upper constraint on $(d_{NN\gamma})$ yields approximately $10^{-2}\,\rm eV^{-1}$ .
 IV.2 Sterile to active neutrino decay
-In this section, we consider the $\mathcal{O}$(eV) mass scale of sterile neutrino, which is decaying into active neutrino and photon via active-to-sterile transition magnetic dipole moment.
-This radiative decay process can be expressed as $N_{1}\to\nu_{a}+\gamma$.
-The wavelength of the emitted photon $(\lambda_{e})$ can be calculated from Eq.
+In this section, we examine light sterile neutrinos in the mass range $m_{s}\in[4,\,20]\,\mathrm{eV}$.
+These neutrinos behave as dark radiation rather than cold dark matter.
+Consequently, their abundance is typically computed using the Dodelson–Widrow (DW) mechanism, given by $\Omega_{s}h^{2}=\frac{m_{s}}{94.1\,\mathrm{eV}}$.
+However, for sterile neutrino masses within this range, the DW mechanism is ruled out by cosmological observations, including data from the Cosmic Microwave Background (CMB) and large-scale structure formation Xu et al.
 (2022).
-The term $\lambda_{e}$ depends on both the mass of sterile and active neutrino.
-Here, we fix the active neutrino mass to be $0.12\,\rm eV$.
-As explained earlier in section (IV.1), the required energy of the photons should be between $2\,\rm eV-10\,\rm eV$ such that these redshifting photons will be observed today as an excess radiation in the COB.
-Using these conditions, we find that the mass of the sterile neutrino will be in the range of $4\,\rm eV-20\,\rm eV$.
-The upper bound on the mass of the sterile neutrino to be a dark matter candidate, derived from the Traemaine-Gunn bound, is $0.4$ keV Abazajian and Koushiappas (2006).
-Therefore, we examine the $\mathcal{O}$(eV) scale decaying sterile neutrinos, which may be considered as dark radiation.
-In the standard scenario case, only three neutrino species contribute to the $\rm N_{eff}^{SM}$ which is $3.044$.
-However, in a nonstandard scenario, the effective neutrino species can be written as $\rm N_{eff}=N_{eff}^{SM}+\Delta N_{eff}$, where $\Delta\rm N_{eff}$ comes from the additional relativistic components that take part in the dark radiation Abazajian and Koushiappas (2006).
-Dark radiation measures the amount of radiation energy contributed by the relativistic species except photons.
-Usually, neutrino oscillation anomalies give the idea for searching the new additional relativistic particles in the cosmos Abazajian and Koushiappas (2006).
-Motivated by the neutrino oscillation experiment for explaining the dark radiation, the mass range of sterile neutrino should be $m_{s}\lesssim\rm eV$ Drewes (2013).
-So here, we take the mass range $(4\,\rm eV-20\,\rm eV)$ of decaying sterile neutrino, which can be regarded as dark radiation.
-Figure 4: (a): Represent the variations in specific intensity $I_{\lambda}$ with respect to $\lambda_{\rm obs}$, for $12$ eV sterile neutrinos decaying at $10^{-23}\,\text{sec}^{-1}$ and $10^{-22}\,\text{sec}^{-1}$, shown in blue solid and red dashed lines, respectively.
-(b): Shows variations in $I_{\lambda}$ for $12$ eV sterile neutrinos of masses with decay rates $10^{-23}-10^{-22}\,\rm sec^{-1}$ for fixed intensities $I_{806}$ and $I_{299}$, as shown in the red dashed and blue solid lines, respectively.
-The green-shaded area illustrates the responsivity curve of LORRI for each detected wavelength.
-The decay width for a decaying sterile neutrino via sterile-to-active transition magnetic moment is given by Beltrán et al. (2024)
-$\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{|d_{\rm\nu N\gamma}|^{2}}{2\pi}m_{1}^{3}\,,$ (9)
-where $d_{\nu\rm N\gamma}$, $m_{1}$, and $m_{2}$ represent the active-to-sterile transition magnetic dipole moment and the mass of the sterile and active neutrino, respectively.
-On the contrary to the decay width shown in Eq.
-(2), here $\Gamma_{m_{1}}$ is independent of the mass difference between the sterile and active neutrino.
-We then write the expression for the mean specific intensity $I_{\lambda}$ of photons produced from eV-scale decaying sterile neutrino as
-$I_{\lambda}=x\,\frac{c}{4\pi}\frac{{\Omega_{s}}{\rho_{c}}{c^{2}}\Gamma_{m_{1}}}{\lambda_{obs}(1+z)H(z)}.$ (10)
-Here, the term $x=1/2\,\left(1-m_{2}^{2}/m_{1}^{2}\right)$ is similar to the one shown in Eq.
+Assuming a conservative minimum relic temperature of $T_{X}=0.91\,\mathrm{K}$, stringent cosmological constraints exist on the mass of thermal light relics.
+Ref.
+Xu et al.
+(2022) analyzes four categories of such relics—real scalars, Weyl fermions, Dirac fermions, and vectors.
+Their analysis shows that minimally coupled Weyl fermions with $m_{X}>2.3\,\mathrm{eV}$, real scalars with $m_{X}>11\,\mathrm{eV}$, and Dirac fermions with $m_{X}>1.1\,\mathrm{eV}$ are excluded at the 95% confidence level Aker et al.
 (2019).
-Whereas, $\Omega_{s}$ represents the density parameter of sterile neutrinos, which can be expressed as $\Omega_{s}h^{2}=m_{1}/94.1\,\rm eV$ for a nonthermal scenario, which comes from the “Dodelson-Widrow Mechanism” Dodelson and Widrow (1994).
-Using Eq.
-(10), we calculate $I_{\lambda}$ for $(4\,\rm eV-20\,\rm eV)$ mass range of sterile neutrinos, that can contribute to the reported intensities by LORRI in the wavelength range of $(0.4\,\rm\mu m-0.9\,\rm\mu m)$.
-For a fixed $m_{1}=12\,\rm eV$, we considered decay widths $\Gamma_{m_{1}}$ of $10^{-23}\,\rm sec^{-1}$ and $-10^{-22}\,\rm sec^{-1}$ to plot variations in $I_{\lambda}$ for different $\lambda_{\rm obs}$ values— shown in blue solid and red dashed lines, respectively, in Fig.
-4.
-Further in Fig.
-4, we vary $m_{1}$ in the range $(4\,\rm eV-20\,\rm eV)$ and $\Gamma_{m_{1}}$ within $10^{-23}-10^{-22}\,\rm sec^{-1}$, while fixing the total intensity of photons to $I_{806}$ and $I_{299}$— as shown in red dashed and blue solid lines, respectively.
-Below, we calculate the decay rate for mass $m_{1}$ required to obtain both the COB intensities and derive upper bounds on sterile-to-sterile transition magnetic dipole moment.
-Figure 5: Top panel: The minimum required decay rate to put constraint on sterile-to-active transition magnetic dipole moment ($d_{\nu N\gamma}$) by taking into consideration the contribution of the anomalous intensity of COB $2.99\pm 2.03\,\rm nW/m^{2}/sr$ Postman et al.
-(2024) and the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al.
-(2022) reported by LORRI for sterile neutrinos with masses ($4\leavevmode\nobreak\ \rm eV-20\leavevmode\nobreak\ \rm eV)$.
-Bottom panel: The corresponding minimum magnetic coupling strength $(d_{\rm\nu N\gamma})$ in $\text{eV}^{-1}$.
-In the top panel of Fig.
-5, the blue solid and red dashed lines show the minimum decay width $\Gamma_{m_{1}}$ for sterile neutrino $4\leavevmode\nobreak\ \text{eV}\leq m_{1}\leq 20\leavevmode\nobreak\ \text{eV}$ to obtain the total intensities of $I_{299}$ and $I_{806}$, respectively.
-We find that $\Gamma_{m_{1}}$ increases for heavier sterile neutrinos.
-This can be analyzed from the fact that the specific intensity $(I_{\lambda})$ depends directly on the decay width while mass enters only through redshift $z(\lambda_{\rm obs},m_{1})$ which can only alter the Hubble parameter $H(z)$— as shown in Eq.
+These bounds are notably more stringent—by factors of 2 to 5—than those reported in previous literature Xu et al.
 (2022).
-Thus, larger mass increases $z(\lambda_{\rm obs},m_{1})$, suggesting an early decay of sterile neutrino.
-However, as the intensity of photons redshifts more for an early decay, a larger decay width is required to obtain the required intensity today.
-Further, the solid blue and dashed red lines in the bottom panel show upper bounds on active-to-sterile transition magnetic moment $d_{\nu N\gamma}$ derived for the above blue solid line and red dashed line respectively using Eq.
-(9).
-We observe that $d_{\nu N\gamma}$ decreases with the increase of the mass of the decaying sterile neutrino.
-This can be analysed from Eq.
-(9), where $d_{\nu N\gamma}$ is inversely proportional to the mass of decaying sterile neutrinos.
-We find that, the required $d_{\nu N\gamma}$ to explain the anomalous intensities of $I_{806}$ and $I_{299}$ lies between $3.33\times 10^{-12}\,\rm\mu_{B}-10^{-11}\,\rm\mu_{B}$ and $3.33\times 10^{-13}\,\rm\mu_{B}-10^{-12}\,\rm\mu_{B}$, respectively, for sterile neutrino mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$.
-Currently, the upper bound on the active-to-sterile transition dipole moment $d_{\rm\nu N\gamma}$ for sterile neutrino’s mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$ remains unconstrained.
-There are several experimental upper bounds on the active-to-sterile transition magnetic dipole moment.
-Recent studies of XENON1T found that the upper bound on $d_{\nu N\gamma}\lesssim 6\times 10^{-11}\mu_{B}$ for several MeV mass range of sterile neutrino Shoemaker et al.
-(2021).
-However, the recent analysis of global clusters sets an upper limit as $d_{\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Cañas et al.
-(2016).
-IV.3 Active-to-active neutrino magnetic moment
-In the UV model, the active-to-active neutrino magnetic moment is naturally suppressed Beltrán et al.
+Furthermore, results from CMB analyses, weak gravitational lensing Abbott et al.
+(2022), and galaxy surveys strongly constrain light thermal relics with masses $m_{X}\gtrsim 0.1\,\mathrm{eV}$.
+Therefore, the observed excess in the Cosmic Optical Background (COB) intensity cannot be attributed to the decay of sterile neutrinos into active neutrinos within the considered mass range of $4\,\mathrm{eV}\leq m_{s}\leq 20\,\mathrm{eV}$.
+IV.3 Active-to-active neutrino transition magnetic moment
+In the UV model, the active-to-active neutrino transition magnetic moment is naturally suppressed Beltrán et al.
 (2024).
@@
 For that reason, we cannot observe any specific intensity of photons in the observed wavelength range $\lambda_{obs}=0.2\leavevmode\nobreak\ \mu m-1.2\leavevmode\nobreak\ \mu m$ because to explain the COB anomaly, we need minimum $4$ eV mass of sterile neutrinos.
-To summarize, we put constraints on the sterile-to-sterile and sterile-to-active transition magnetic dipole moment for ($1\,\rm keV-20\,\rm keV$) and ($4\,\rm eV-20\,\rm eV$) mass range of stile neutrinos, respectively by considering two COB measurements.
-V Conclusion and Outlook
+Hence, the COB anomalous intensities can not be accounted for by the decay of active neutrinos into active neutrinos and photons.
+V UV Origin of Small Mass Splittings in keV Sterile Neutrinos
+Although our primary analysis adopts an effective field theory approach, a complete understanding of the small eV-scale mass splitting between keV sterile neutrinos benefits from a consistent ultraviolet (UV) completion.
+Several well-motivated frameworks exist that naturally explain the emergence of quasi-degenerate sterile neutrino masses without fine-tuning.
+Two particularly appealing UV approaches are as follows:
+  1. 1.
+    Froggatt–Nielsen Mechanism: In models with a $U(1)_{\rm FN}$ Froggatt–Nielsen symmetry Froggatt and Nielsen (1979); Pilaftsis and Underwood (2004), sterile neutrinos are assigned flavor-dependent charges under $U(1)_{\rm FN}$, and their Majorana masses arise from non-renormalizable operators suppressed by a high scale $\Lambda_{\rm FN}$ (the Froggatt–Nielsen (FN) cutoff scale, typically taken to be the GUT or Planck scale).
+The FN symmetry constrains the structure of the mass matrix, enforcing degeneracy at leading order.
+When the scalar field $S$ acquires a vacuum expectation value (VEV), the symmetry is spontaneously broken, generating mass terms suppressed by powers of the small parameter $\epsilon=\langle S\rangle/\Lambda_{\rm FN}$.
+Subleading contributions to the mass splitting may also arise from higher-order corrections or small perturbations consistent with the FN charge assignments.
+These effects induce naturally small mass splittings of the form
+    $\Delta m\sim\epsilon^{n}\Lambda_{\rm FN}$
+    where $n$ depends on the FN charge assignment and operator dimension.
+    In particular, if two sterile neutrinos $N_{1}$ and $N_{2}$ are assigned suitable FN charges, they can acquire masses at the keV scale via higher-dimensional operators suppressed by $\epsilon=\langle S\rangle/M_{\rm GUT}$.
+The resulting Majorana mass terms exhibit the structure:
+    $m_{N_{1,2}}\sim\epsilon M\sim\text{keV},\quad\Delta m_{12}\sim\epsilon^{2}M\sim\text{eV}$
+    for $M\sim\text{TeV}$, where $M$ denotes the mass of a heavy mediator fermion $X$ that appears in the UV completion and is integrated out to generate the effective operator and $\epsilon\sim 10^{-6}$.
+This setup yields quasi-degenerate keV-scale sterile neutrinos with eV-level splittings—without requiring fine-tuning—making them viable dark matter candidates.
+The underlying process is diagrammatically shown in Figure 4.
+[…]
+    Figure 4: Majorana mass generation for sterile neutrino $N$ via two insertions of the FN scalar $S$, whose VEV breaks the $U(1)_{\rm FN}$ symmetry.
+The interaction is mediated by a heavy fermion $X$ of mass $M$.
+  2. 2.
+    Clockwork or Radiative Symmetry Breaking: In alternative frameworks, the degeneracy of sterile neutrino masses is protected by a discrete or continuous symmetry that is broken radiatively.
+This results in small loop-induced splittings between otherwise degenerate states.
+The clockwork mechanism, in particular, achieves exponential suppression of mass terms through a chain of fields with nearest-neighbor interactions Giudice and McCullough (2017); Maharana and Srivastava (2025); Hong et al.
+(2019); Hambye et al.
+(2004); Blanchet et al.
+(2010).
+When applied to sterile neutrinos, such setups can yield keV-scale mass eigenstates with eV-level splittings, making them phenomenologically viable and technically natural.
+These UV completions demonstrate that small splittings among sterile neutrinos can emerge from structural features of the theory, rather than ad hoc fine-tuning, and are compatible with broader frameworks like seesaw and extra-dimensional models.
+Thus, the small $\mathcal{O}(\text{eV})$ mass difference between keV sterile states—essential for oscillation and decay phenomenology—can be realized in a variety of UV-consistent ways.
+VI Conclusion and Outlook
 Recent measurements of the cosmic optical background (COB) photon flux intensity by LORRI have revealed some excess intensity with no known sources.
 This excess may indicate a potential source of energetic photons beyond the Standard Model (BSM).
-In this study, we considered the possibility of decaying sterile neutrinos producing photons and explored the corresponding decay rates.
+In this study, we investigated the possibility of sterile neutrinos decaying into photons and explored the corresponding decay widths.
 Notably, we did not account for photon energy generation through cascades involving other decay products.
-We found the required decay rate of sterile neutrino and obtained upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments of the order $\mathcal{O}(10^{-22}-10^{-21})\leavevmode\nobreak\ \text{sec}^{-1}$ for the keV mass range of the decaying sterile neutrino and for the $\mathcal{O}$(eV) mass range of the decaying sterile neutrino, the decay rate should be in the range of $\mathcal{O}(10^{-23}-10^{-22})\leavevmode\nobreak\ \text{sec}^{-1}$.
+We found the required decay width of the sterile neutrino and obtained upper bounds on the sterile-to-sterile transition magnetic moments of the order $\mathcal{O}(10^{-22}-10^{-21})\leavevmode\nobreak\ \text{sec}^{-1}$ for the keV mass range of the decaying sterile neutrino.
 Here, we adopt a minimal framework involving sterile neutrinos using the effective field theory approach.
-The radiative decay of sterile neutrinos, resulting in photon production, is primarily governed by magnetic moments associated with sterile-to-sterile $(d_{NN\gamma})$ and sterile-to-active neutrino ($d_{\nu N\gamma}$) transitions.
-We calculate the $d_{NN\gamma}$ and $d_{\nu N\gamma}$ for various mass scales and decay rates.
-We find that, $d_{NN\gamma}$ take values of $3\times 10^{-11}\,\mu_{B}-10^{-9}\,\mu_{B}$ for sterile neutrinos of mass range $1\,\rm keV-40\,\rm keV$.
-Whereas sterile neutrinos, with mass range $4\,\rm eV-20\,\rm eV$, decaying via sterile-to-active transition magnetic moment $(d_{\nu\rm N\gamma})$ take values of $3.33\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$ for newly found COB measurement.
-In our work, we explain the bounds on sterile-to-sterile and sterile-to-active transition magnetic moments by considering the two COB measurements: one is $2.99\pm 2.03\,\rm nW/m^{2}/sr$ intensity, and another one is due to anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$.
+The radiative decay of sterile neutrinos, resulting in photon production, is primarily governed by magnetic moments associated with the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$.
+We calculate the $d_{NN\gamma}$ for the different $\Delta m$ values, which vary from $2-10\,\rm eV$.
+We find that, $d_{NN\gamma}$ take values of $3\times 10^{-12}\,\rm eV^{-1}-10^{-9}\,\rm eV^{-1}$ for sterile neutrinos of mass range $1\,\rm keV-40\,\rm keV$.
+In our work, we explain the bounds on sterile-to-sterile transition magnetic moments by considering the two COB measurements: one is $2.99\pm 2.03\,\rm nW/m^{2}/sr$ intensity, and the other one is due to anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$.
 Here, we have plotted the specific intensity for both COB measurements.
-Motivated by this work, one can probe CUB, CXB, and CIB measurements to constrain the sterile-to-sterile and sterile-to-active transition magnetic dipole moment.
+Motivated by this work, one can probe CUB, CXB, and CIB measurements to constrain the sterile-to-sterile transition magnetic moment.
 We leave this as future work.
 Some upcoming experiments, such as GALEX and ULTRASAT satellites, can probe and constrain the decaying dark matter in the ultraviolet bandwidth Libanore and Kovetz (2025).
-Our minimal framework based on decaying sterile neutrinos of mass $\mathcal{O}$(keV) and $\mathcal{O}$(eV) can contribute to the unknown excess COB intensity, thus deriving upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments.
-VI Acknowledgement
+Our minimal framework based on decaying sterile neutrinos of mass $\mathcal{O}$(keV) can contribute to the unknown excess COB intensity, thus deriving upper bounds on the sterile-to-sterile transition magnetic moments.
+In conclusion, our analysis establishes that the anomalous rest-frame intensity of the Cosmic Optical Background enables constraints on the sterile-to-sterile transition magnetic moment, $d_{NN\gamma}$, while scenarios involving sterile-to-active or active-to-active transitions remain disfavored within this framework.
+VII Acknowledgement
 We sincerely acknowledge the usage of python packages scipy [^1] https://www.scipy.org/ and astropy [^2] https://www.astropy.org/index.html.
 We thank Rahul Kothari for his earlier collaboration on this project.
+H.
+H.
+acknowledges support from the Ministry of Higher Education, Government of India.
+T.
+S.
+is supported by the National Natural Science Foundation of China under Grants No.
+12475094, No.
+12135006, and No.
+12075097.